OpenStudy (anonymous):

cos 1395.find d value?

OpenStudy (anonymous):

what does d represent here?

OpenStudy (anonymous):

-1/root2

OpenStudy (anonymous):

hey shruti do u remember me... u wr asking me abt complex nos once?

OpenStudy (sriram):

cosine of any number repeats after every 360 degrees therefore for getting the answer keep subtracting 360 till u r getting a +ve result here 1395-360=1035 1035-360=675 675-360=315 315-360=-45 all the angles on the right of '=' sign have same values of trigo funtions hence cos1395= cos(-45) if u know cos(-x)=cos(x) therefore the answer is 1/square root2